OpenStudy (anonymous):

Find the area of three-fifths of a circle with radius 15 ft. A. 89 B. 283 C. 424 D. 524

OpenStudy (anonymous):

(3/5) * pi* 15^2

OpenStudy (anonymous):

(3/5)A = (3/5) pi * r^2 -> (3/5) pi * 255 = 424 ft^2